A crew can row 8km downstreat and back in 4.8 hours. If the rate of the current is 4 kph, find the rate of the crew in still water.

Oh please help me!!
Matt

Answer by venugopalramana(3286)

A crew can row 8km downstreat and back in 4.8 hours. If the rate of the current is 4 kph, find the rate of the crew in still water.
let crew's speed in still water =v kph
water current = 4 kph
speed up stream = v-4 kph
speed down stream =v+4 kph
total time to ride 8 km up and down=8{1/(v-4)+1/(v+4)=16v/(v^2-16)=4.8
4.8v^2-16*4.8=16v
0.3 v^2-v-4.8=0
3v^2-10v-48=0
3v^2-18v+8v-48=0
3v(v-6)+8(v-6)=0
(v-6)(3v+8)=0
v=6